CARBON IN THE FORM OF GRAPHITE COMBINES WITH EXCESS HYDROGEN GAS TO FORM BENZENE C6H6 in the following reaction 3.95kj OF HEAT WERE TRANSFERRED. Calculate the grams of graphite reacted. Is the reaction endothermic or exothermic
Answer #1

The balanced reaction is

6 C ( graphite, s) + 3 H2 (g) C6H6 (l)

Now, standard heat of formation of Benzene = + 49.2 KJ/mol.

In the reaction heat is absorbed hence, it is endothermic.

So, moles of Benzene firmed when 3.95 KJ of heat is transferred

= 3.95 (KJ) ×

= 0.080 .

Now, from the balanced equation mole ratio of Graphite and Benzene = 6 : 1.

Therefore, moles of graphite reacted

= Moles of benzene × 6

= (0.080 × 6)

= 0.48.

Now, mass of graphite needed

= Moles of graphite × atomic mass of graphite (carbon)

= 0.48 (mol) × 12 (g/mol)

= 5.78 g